OpenStudy (anonymous):

Find all values of x satisfying the Mean Value Theorem for f(x) = x^3 on [0, 1].

OpenStudy (anonymous):

Ok first we use the mean value theorem equation \[f(c)=\frac{ 1 }{ b-a }\int\limits_{a}^{b}f(x) dx\]

OpenStudy (anonymous):

so b=1 and a=0

OpenStudy (anonymous):

and f(x)=x^3

OpenStudy (anonymous):

Plug in those things into the equation

OpenStudy (anonymous):

Then tell me the result